OpenStudy (anonymous):

write an exponential function whose graph passes through points (5,729) and (3,81)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Well it has a base of 3

OpenStudy (anonymous):

ok how did you get that equation

OpenStudy (luigi0210):

Lucky guessing

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you can actually compute it without guessing if you like

OpenStudy (anonymous):

you are told \((3,81)\) is on the graph of \(y=ab^x\) which means \(81=ab^3\) similarly you know \(729=ab^5\) divide the second by the first

OpenStudy (anonymous):

this tells you \[\frac{729}{81}=\frac{ab^5}{ab^3}\] or \[9=b^2\] making \(b=3\) guessing is good too!

OpenStudy (anonymous):

really how... because if I can get the idea of how to do one I may be able to DO the other problems I have by myself.

OpenStudy (anonymous):

then since \(81=a\times 3^3=27a\) you know \(a=3\) as well

OpenStudy (luigi0210):

haha, satellite's a genius

OpenStudy (anonymous):

making the function \[f(x)=3\times 3^x\] or more commonly written as \[f(x)=3^{x+1}\]

OpenStudy (luigi0210):

my equation again: \[y=3^{x+1}\]

OpenStudy (anonymous):

like i said, guessing is a good method as well

OpenStudy (luigi0210):

And sorry ML, once you get enough practice these will be second nature to you

OpenStudy (anonymous):

OK thank YOU guys I will go on & try to do the others with this blueprint